Abigail is three times as old as Ryan. In 6 years time Abigail will be 42. How old will Ryan be in two years?
Answer by FrankM(1040)

In 6 years time Abigail will be 42
Abigail is 42-6= 36
Abigail is three times as old as Ryan
Ryan is 36/3 = 12 now
How old will Ryan be in two years?
12+2=14